# Almost real closed fields with real analytic structure

### v1

@@ 3. Analysing terms
Let ${\mathcal{L}}_{\mathcal{B}}$ be the language $\{0,1,+,\cdot,(\cdot)^{-1},<\}$ together with function symbols for all elements of ${\mathcal{B}}$. Then the field $K$ is naturally an ${\mathcal{L}}_{\mathcal{B}}$-structure. Even though we will consider $K$ as a valued field for the canonical valuation $\lVert\cdot\rVert:K^{\times}\to G$, note that ${\mathcal{L}}_{\mathcal{B}}$ does not have a symbol for the valuation ring.
@@ 3.1. An embedding theorem
Let $K$ be an almost real closed field, with canonical valuation $\lVert\cdot\rVert\colon K^{\times}\to G$. Then, for any subgroup $\Gamma\leq G$, any partial section $s_{0}\colon\Gamma\to K_{>0}$ of the valuation extends to a section $s\colon G\to K_{>0}$.
@@
By Lemma 2.1.10 there is a unique extension of the ${\mathcal{B}}$-analytic structure on $L$ and $K$ to their real closures $L^{\mathrm{rcl}}$ and $K^{\mathrm{rcl}}$. Inside $K^{\mathrm{rcl}}$, the resulting $\mathrm{qftp}_{{\mathcal{L}}_{{\mathcal{B}}}}(a/L^{\mathrm{rcl}})$ clearly contains $\mathrm{qftp}_{{\mathcal{L}}_{{\mathcal{B}}}}(a/L)$. Since the ${\mathcal{L}}_{{\mathcal{B}}}$-structure on $L^{\mathrm{rcl}}$ is o-minimal by [5, Thm. 3.4.3], it follows that $\mathrm{qftp}_{{\mathcal{L}}_{{\mathcal{B}}}}(a/L^{\mathrm{rcl}})$ is determined by $\mathrm{qftp}_{\{<\}}(a/L^{\mathrm{rcl}})$. By uniqueness of the real closure, the latter is completely determined by $\mathrm{qftp}_{{\mathcal{L}}_{\mathrm{or}}}(a/L)$. ∎
The following embedding theorem can be seen as an extension of an embedding result by van den Dries–Macintyre–Marker [25, Cor. 3.6], and is the main result of this section. Recall that if $\Omega$ is a a subgroup of an ordered abelian group $G$ and ${\mathcal{B}}$ is a Weierstrass system over $\Omega$, then ${\mathbb{R}}((G))$ has a natural ${\mathcal{L}}_{{\mathcal{B}}}$-structure, as described in Example 2.1.4.
@@
Let ${\mathcal{B}}\subset{\mathcal{A}}((\Omega))$ be a rich Weierstrass system and let $K$ be an ordered field with ${\mathcal{B}}$-analytic structure, with value group $G$ for the canonical valuation. Then there exists an inclusion $\Omega\hookrightarrow G$ of abelian groups, and an ${\mathcal{L}}_{{\mathcal{B}}}$-embedding $K\hookrightarrow{\mathbb{R}}((G))$ where ${\mathbb{R}}((G))$ is equipped with the natural ${\mathcal{B}}$-analytic structure.
@@
The inclusions ${\mathbb{R}}(\Omega)\subset B_{0}\subset K$ determine an inclusion $\Omega\leq G$ and a partial section $s\colon\Omega\to B_{0}\subset K_{>0}\colon\omega\to t^{\omega}$ for the canonical valuation on $K$. By Lemma 3.1.1, $s$ extends to a section $G\to K_{>0}$. Additionally the inclusion $\Omega\leq G$ induces an ${\mathcal{L}}_{{\mathcal{B}}}$-embedding $\iota_{0}\colon B_{0}\hookrightarrow{\mathbb{R}}((G))$. Using Zorn’s lemma, we construct a maximal ${\mathcal{L}}_{{\mathcal{B}}}$-substructure $L$ of $K$ together with an ${\mathcal{L}}_{{\mathcal{B}}}$-embedding $\iota\colon L\hookrightarrow{\mathbb{R}}((G))$, extending the one on $B_{0}.$ We show that $L=K$. Assume that $L\neq K$ and consider the following cases, where we write $\Gamma_{L}$ for the value group of $L$ with respect to the canonical valuation:
@@
(3) Since $L\neq K$, there exists some $\lambda\in{K}^{\circ\circ}\setminus{L}^{\circ\circ}$. Because $K$ is an immediate extension of $L$, we may find a sequence $(x_{\alpha})_{\alpha<\beta}$ in $L$ which pseudo-converges to $\lambda$ and which has no pseudo-limit in $L$ [13, p. 104]. Now let $\mu\in{\mathbb{R}}((G))$ be any a pseudo-limit of $(x_{\alpha})_{\alpha<\beta}$ in ${\mathbb{R}}((G))$, which cannot be an element of $L$. As in case (2), it suffices to show that for any polynomial $p(x)\in L[x]$ we have $p(\lambda)>0$ if and only if $p(\mu)>0$. Now notice that eventually $\lVert p(x_{\alpha})-p(\lambda)\rVert<\lVert p(\lambda)\rVert$, since the left hand side is eventually strictly decreasing and $\lVert p(x_{\alpha})\rVert$ is eventually constant. Indeed, $(x_{\alpha})_{\alpha<\beta}$ is of transcendental type, as $L$ is henselian of equicharacteristic zero, by Lemmas 2.1.9 and [13, Corollary 4.22]. In particular, for all sufficiently large $\alpha$ we have $\lvert p(x_{\alpha})-p(\lambda)\rvert<\lvert p(\lambda)\rvert$. Hence the sign of $p(x_{\alpha})$ is eventually constant, and equal to the sign of $p(\lambda)$. The same holds for $p(\mu)$, whence it has the same sign as $p(\lambda)$. ∎
@@ 3.2. Extending real Weierstrass systems
Let $A$ be a subset of $K$. The goal of this section is to extend the Weierstrass system ${\mathcal{B}}$ with elements from $A$, similar to [5, Sec. 4.5] for separated Weierstrass systems. The following proposition is the main goal.
@@
Let ${\mathcal{B}}$ be a real Weierstrass system over $\Omega\leq G$. Given an ${\mathcal{L}}_{\mathcal{B}}$-substructure $A$ of ${\mathbb{R}}((G))$, we define a family of ${\mathbb{R}}$-algebras ${\mathcal{B}}[A]=\{B_{m,\alpha}[A]\}_{m,\alpha}$ where $n$ ranges over ${\mathbb{N}}$ and $\alpha$ over ${\mathbb{R}}_{>0}$:
@@
By the Strong Noetherian Property [5, Thm. 3.2.2], we may write
@@ 3.3. Rings of analytic functions
The requirement that ${\mathcal{B}}$ is strong, guarantees that ${\mathcal{O}}_{D}({\mathcal{B}})$ is, in fact, a ring. For more details, see [5] and [8, Remark 3.1].
@@ 3.4. Term analysis for real closed fields
The below proposition is an extension of [8, Prop. 3.13]. The main difference is that we need control on the parameters from $A$. This requires functions from ${\mathcal{B}}[A]$ to be used in the definition of the ${\mathcal{O}}_{X_{i}}$ rather than only functions from ${\mathcal{B}}$.
@@
In [8, Prop. 3.13] it is shown that this proposition holds when $A={\mathbb{R}}((G))$. Hence, there exist $X_{i}$, $\tau_{i}$ as in the statement but with $X_{i}$ defined over ${\mathbb{R}}((G))$ and $\tau_{i}\in{\mathcal{O}}_{X_{i}}({\mathbb{R}}((G)))$. Now this proposition essentially follows from the fact that $A\preccurlyeq{\mathbb{R}}((G))$ as ${\mathcal{L}}_{{\mathcal{B}}[A]}$-structures. We give some details.
@@
Now recall that we have quantifier elimination in ${\mathcal{L}}_{\mathcal{B}}$ for the theory of real closed fields with ${\mathcal{B}}$-analytic structure, by [5, Thm. 3.4.3]. Hence, $A\preccurlyeq{\mathbb{R}}((G))$ and there exists a covering of $[-1,1]_{A}$ by finitely many intervals and annuli with centers and radii in $A$, such that on each annulus $\tau(x)$ is of the form (3.4.2) and on each interval $\tau(x)$ is of the form (3.4.1) for $a,b\in A$ (except at a finite number of points). If all $a,b$ appearing in (3.4.2) and (3.4.1) can be taken infinitesimal, then the right-hand sides belong to ${\mathcal{O}}_{I}({\mathcal{B}}[A])$ and ${\mathcal{O}}_{D}({\mathcal{B}}[A])$ respectively, and we are done. Suppose this is not the case for some annulus $D$. Then apply axiom (1c) and compactness of $[-1,1]_{{\mathbb{R}}}$ to cut up $D$ into finitely many annuli $D_{1},\dots,D_{q}$ such that $\tau_{|D}$ belongs to ${\mathcal{O}}_{D_{j}}({\mathcal{B}}[A])$ for every $j=1,\dots,q$, except at finitely many points. The case of an interval is similar.
@@
This yields a covering of $[-1,1]_{K}$ by finitely many intervals and annuli $X_{i}$ such that on each of them $\tau(x)$ agrees with an element of ${\mathcal{O}}_{X_{i}}({\mathcal{B}}[A])$, except at finitely many points. Fix any interval or annulus $X$ of this cover of $[-1,1]_{K}$. If $X$ is an interval, then similarly to [8, Cor. 3.5] we may find finitely many intervals $\{I_{i}\}_{i=1}^{n}$ covering $X$, polynomials $P_{i}(x)\in K[x]$ and units $U_{i}(x)\in{\mathcal{O}}_{I_{i}}^{\times}$ such that $\tau_{|I_{i}}(x)=P_{i}(x)U_{i}(x)$ except at finitely many points, for all $i=1,\dots,n$. Then use Lemma 3.3.5 to control $\operatorname{rv}_{\lambda}(U_{i}(x))$.
If $X=D(c,\delta,\varepsilon)$ is an annulus, then apply the same arguments as in [8, Lemma 3.7]. We may split of any number of intervals, since we already know how to deal with these. So let $D=D(c,\delta,\varepsilon)$ be one of the finitely many resulting annuli. The arguments of [8, Lemma 3.7] show that
@@ 3.5. Pulling back from the real closure
We use induction on $n=\deg(p)$. If $n=1$ it is clear, since then we can simply take $C=\{\alpha_{1}\}$. So we assume that $n>1$ and that the statement is proven for smaller $n$. By induction, let $C$ be a finite set for the claim applied to $p^{\prime}$, and also such that $\operatorname{rv}(p(x))$ is constant on balls $1$-next to $C$, which is possible by [14, Prop. 3.6]. This $C$ is ${\mathcal{L}}_{\mathrm{val}}(A)$-definable, and we claim that it suffices. So let $x,y$ be in the same ball $1$-next to $C$, and assume that the claim is false. Then there exists some root $\alpha$ of $p$ such that $\operatorname{rv}(x-\alpha)\neq\operatorname{rv}(x-\alpha)$. Let $B\subset K^{\mathrm{rcl}}$ be the closed valuative ball around $x$ of radius $v(x-y)$, then $\alpha$ is an element of $B$. If $B$ would contain another root $\alpha^{\prime}\in K^{\mathrm{rcl}}$ of $p$, then by Rolle’s theorem it would also contain a root of $p^{\prime}$, since $K^{\mathrm{rcl}}$ is real closed and $B$ is convex. But this contradicts the construction of $C$, and so the only root of $p$ in $B$ is $\alpha$. Hence if $\alpha^{\prime}$ is another root of $p$, then $\operatorname{rv}(x-\alpha^{\prime})=\operatorname{rv}(y-\alpha^{\prime})$. But if we now compute
@@ 4.1. Eliminating to $\mathrm{RV}_{\lambda}$
The proof of h-minimality is based on relative quantifier elimination results in a larger many-sorted language. This quantifier elimination is similar to work by Denef–van den Dries [10], Pas [24] and Flenner [14]. To go from relative quantifier elimination to $\omega$-h-minimality, we follow the method from [3, Sec. 6].
@@
We call this the partial addition. We recall a first relative quantifier elimination result due to Flenner.
4.1.1 Theorem ([14, Prop. 4.3]).
@@
where $z=(z_{1},\ldots,z_{m})$ are valued field variables, the $f_{i}$ are either in $B_{m+n,\alpha}$ or a polynomial over $K$, and no field division occurs. Let $(z_{0},y_{0})\in[-1,1]^{m+n}$. Using the Strong Noetherian property from [5, Theorem 3.2.2], there exist an integer $d$ and a real number $\beta(z_{0},y_{0})>0$, unit elements $u_{ij}\in B^{\circ}_{m+n,\beta}$, elements $h_{ij}\in B_{n,\beta}$ and a subset $J_{i}$ of $\{0,...,d\}^{m}$ such that for every $i=1,\ldots,s$,
@@
for finitely many points $(z_{0},y_{0})\in[-1,1]_{{\mathbb{R}}}^{m+n}$. Without loss of generality, we can assume that $(z_{0},y_{0})=(0,0)$. Now, by a coordinate transformation and Weierstrass division similar to [8, Proof of Theorem 2.14] and [10, Lemma 4.13 and Proof 4.14], we can shrink $\beta$ if needed and write $f_{i}(z,y)$ as
@@
for some fixed $j$, $c\in[-1,1]_{{\mathbb{R}}}$, where $v_{i}$ is a unit and $g_{i}$ polynomial in the last variable $z_{m}$. Now, if $v\in B_{n^{\prime},\alpha}^{\circ}$ is a unit, then by construction of our language, there exists a definable map $\overline{v}$ on $\mathrm{RV}_{\lambda}^{n^{\prime}}$ such that $\operatorname{rv}_{\lambda}(v(x))=\overline{v}(\operatorname{rv}_{\lambda}(x))$. Hence $\operatorname{rv}_{\lambda}(f_{i}(z,y))$ depends only on $\operatorname{rv}_{\lambda}(z,y),\operatorname{rv}_{\lambda}(h_{i\ell}(y)/h_{j\ell}(y)-c)$ and $\operatorname{rv}_{\lambda}(g_{i}(y,h_{i\ell}(y)/h_{j\ell}(y)-c,z))$. In other words, the variable $z_{m}$ only occurs polynomially in our formula (with some parameters depending on $y$). If $\lambda=1$, we can then conclude from Theorem 4.1.1 to eliminate the quantifier over $z_{m}$ and iterate this procedure. If $\lambda\neq 1$, then the result follows from [3, Prop. 6.2.3]. ∎

### v2

@@ 3. Analysing terms
Let ${\mathcal{L}}_{\mathcal{B}}$ be the language $\{0,1,+,\cdot,(\cdot)^{-1},<\}$ together with function symbols for all elements of ${\mathcal{B}}$. Then the field $K$ is naturally an ${\mathcal{L}}_{\mathcal{B}}$-structure. Even though we will consider $K$ as a valued field for the natural valuation $\lVert\cdot\rVert:K^{\times}\to G$, note that ${\mathcal{L}}_{\mathcal{B}}$ does not have a symbol for the valuation ring.
@@ 3.1. An embedding theorem
Let $K$ be an almost real closed field, with natural valuation $\lVert\cdot\rVert\colon K^{\times}\to G$. Then, for any subgroup $\Gamma\leq G$, any partial section $s_{0}\colon\Gamma\to K_{>0}$ of the valuation extends to a section $s\colon G\to K_{>0}$.
@@
By Lemma 2.1.10 there is a unique extension of the ${\mathcal{B}}$-analytic structure on $L$ and $K$ to their real closures $L^{\mathrm{rcl}}$ and $K^{\mathrm{rcl}}$. Inside $K^{\mathrm{rcl}}$, the resulting $\mathrm{qftp}_{{\mathcal{L}}_{{\mathcal{B}}}}(a/L^{\mathrm{rcl}})$ clearly contains $\mathrm{qftp}_{{\mathcal{L}}_{{\mathcal{B}}}}(a/L)$. Since the ${\mathcal{L}}_{{\mathcal{B}}}$-structure on $L^{\mathrm{rcl}}$ is o-minimal by [8, Thm. 3.4.3], it follows that $\mathrm{qftp}_{{\mathcal{L}}_{{\mathcal{B}}}}(a/L^{\mathrm{rcl}})$ is determined by $\mathrm{qftp}_{\{<\}}(a/L^{\mathrm{rcl}})$. By uniqueness of the real closure, the latter is completely determined by $\mathrm{qftp}_{{\mathcal{L}}_{\mathrm{or}}}(a/L)$. ∎
The following embedding theorem can be seen as an extension of an embedding result by van den Dries–Macintyre–Marker [18, Cor. 3.6], and is the main result of this section. Recall that if $\Omega$ is a a subgroup of an ordered abelian group $G$ and ${\mathcal{B}}$ is a Weierstrass system over $\Omega$, then ${\mathbb{R}}((G))$ has a natural ${\mathcal{L}}_{{\mathcal{B}}}$-structure, as described in Example 2.1.4.
@@
Let ${\mathcal{B}}\subset{\mathcal{A}}((\Omega))$ be a rich Weierstrass system and let $K$ be an ordered field with ${\mathcal{B}}$-analytic structure, with value group $G$ for the natural valuation. Then there exists an inclusion $\Omega\hookrightarrow G$ of abelian groups, and an ${\mathcal{L}}_{{\mathcal{B}}}$-embedding $K\hookrightarrow{\mathbb{R}}((G))$ where ${\mathbb{R}}((G))$ is equipped with the natural ${\mathcal{B}}$-analytic structure.
@@
The inclusions ${\mathbb{R}}(\Omega)\subset B_{0}\subset K$ determine an inclusion $\Omega\leq G$ and a partial section $s\colon\Omega\to B_{0}\subset K_{>0}\colon\omega\to t^{\omega}$ for the natural valuation on $K$. By Lemma 3.1.1, $s$ extends to a section $G\to K_{>0}$. Additionally the inclusion $\Omega\leq G$ induces an ${\mathcal{L}}_{{\mathcal{B}}}$-embedding $\iota_{0}\colon B_{0}\hookrightarrow{\mathbb{R}}((G))$. Using Zorn’s lemma, we construct a maximal ${\mathcal{L}}_{{\mathcal{B}}}$-substructure $L$ of $K$ together with an ${\mathcal{L}}_{{\mathcal{B}}}$-embedding $\iota\colon L\hookrightarrow{\mathbb{R}}((G))$, extending the one on $B_{0}.$ We show that $L=K$. Assume that $L\neq K$ and consider the following cases, where we write $\Gamma_{L}$ for the value group of $L$ with respect to the natural valuation:
@@
(3) Since $L\neq K$, there exists some $\lambda\in{K}^{\circ\circ}\setminus{L}^{\circ\circ}$. Because $K$ is an immediate extension of $L$, we may find a sequence $(x_{\alpha})_{\alpha<\beta}$ in $L$ which pseudo-converges to $\lambda$ and which has no pseudo-limit in $L$ [17, p. 104]. Now let $\mu\in{\mathbb{R}}((G))$ be any a pseudo-limit of $(x_{\alpha})_{\alpha<\beta}$ in ${\mathbb{R}}((G))$, which cannot be an element of $L$. As in case (2), it suffices to show that for any polynomial $p(x)\in L[x]$ we have $p(\lambda)>0$ if and only if $p(\mu)>0$. Now notice that eventually $\lVert p(x_{\alpha})-p(\lambda)\rVert<\lVert p(\lambda)\rVert$, since the left hand side is eventually strictly decreasing and $\lVert p(x_{\alpha})\rVert$ is eventually constant. Indeed, $(x_{\alpha})_{\alpha<\beta}$ is of transcendental type, as $L$ is henselian of equicharacteristic zero, by Lemmas 2.1.9 and [17, Corollary 4.22]. In particular, for all sufficiently large $\alpha$ we have $\lvert p(x_{\alpha})-p(\lambda)\rvert<\lvert p(\lambda)\rvert$. Hence the sign of $p(x_{\alpha})$ is eventually constant, and equal to the sign of $p(\lambda)$. The same holds for $p(\mu)$, whence it has the same sign as $p(\lambda)$. ∎
@@ 3.2. Extending real Weierstrass systems
Let $A$ be a subset of $K$. The goal of this section is to extend the Weierstrass system ${\mathcal{B}}$ with elements from $A$, similar to [8, Sec. 4.5] for separated Weierstrass systems. The following proposition is the main goal.
@@
Let ${\mathcal{B}}$ be a real Weierstrass system over $\Omega\leq G$. Given an ${\mathcal{L}}_{\mathcal{B}}$-substructure $A$ of ${\mathbb{R}}((G))$, we define a family of ${\mathbb{R}}$-algebras ${\mathcal{B}}[A]=\{B_{m,\alpha}[A]\}_{m,\alpha}$ where $m$ ranges over ${\mathbb{N}}$ and $\alpha$ over ${\mathbb{R}}_{>0}$:
@@
By the Strong Noetherian Property [8, Thm. 3.2.2], we may write
@@ 3.3. Rings of analytic functions
The requirement that ${\mathcal{B}}$ is strong, guarantees that ${\mathcal{O}}_{D}({\mathcal{B}})$ is, in fact, a ring. For more details, see [8] and [11, Remark 3.1].
@@ 3.4. Term analysis for real closed fields
The below proposition is an extension of [11, Prop. 3.13]. The main difference is that we need control on the parameters from $A$. This requires functions from ${\mathcal{B}}[A]$ to be used in the definition of the ${\mathcal{O}}_{X_{i}}$ rather than only functions from ${\mathcal{B}}$.
@@
In [11, Prop. 3.13] it is shown that this proposition holds when $A={\mathbb{R}}((G))$. Hence, there exist $X_{i}$, $\tau_{i}$ as in the statement but with $X_{i}$ defined over ${\mathbb{R}}((G))$ and $\tau_{i}\in{\mathcal{O}}_{X_{i}}({\mathbb{R}}((G)))$. Now this proposition essentially follows from the fact that $A\preccurlyeq{\mathbb{R}}((G))$ as ${\mathcal{L}}_{{\mathcal{B}}[A]}$-structures. We give some details.
@@
Now recall that we have quantifier elimination in ${\mathcal{L}}_{\mathcal{B}}$ for the theory of real closed fields with ${\mathcal{B}}$-analytic structure, by [8, Thm. 3.4.3]. Hence, $A\preccurlyeq{\mathbb{R}}((G))$ and there exists a covering of $[-1,1]_{A}$ by finitely many intervals and annuli with centers and radii in $A$, such that on each annulus $\tau(x)$ is of the form (3.4.2) and on each interval $\tau(x)$ is of the form (3.4.1) for $a,b\in A$ (except at a finite number of points). If all $a,b$ appearing in (3.4.2) and (3.4.1) can be taken infinitesimal, then the right-hand sides belong to ${\mathcal{O}}_{I}({\mathcal{B}}[A])$ and ${\mathcal{O}}_{D}({\mathcal{B}}[A])$ respectively, and we are done. Suppose this is not the case for some annulus $D$. Then apply axiom (1c) and compactness of $[-1,1]_{{\mathbb{R}}}$ to cut up $D$ into finitely many annuli $D_{1},\dots,D_{q}$ such that $\tau_{|D}$ belongs to ${\mathcal{O}}_{D_{j}}({\mathcal{B}}[A])$ for every $j=1,\dots,q$, except at finitely many points. The case of an interval is similar.
@@
This yields a covering of $[-1,1]_{K}$ by finitely many intervals and annuli $X_{i}$ such that on each of them $\tau(x)$ agrees with an element of ${\mathcal{O}}_{X_{i}}({\mathcal{B}}[A])$, except at finitely many points. Fix any interval or annulus $X$ of this cover of $[-1,1]_{K}$. If $X$ is an interval, then similarly to [11, Cor. 3.5] we may find finitely many intervals $\{I_{i}\}_{i=1}^{n}$ covering $X$, polynomials $P_{i}(x)\in K[x]$ and units $U_{i}(x)\in{\mathcal{O}}_{I_{i}}^{\times}$ such that $\tau_{|I_{i}}(x)=P_{i}(x)U_{i}(x)$ except at finitely many points, for all $i=1,\dots,n$. Then use Lemma 3.3.5 to control $\operatorname{rv}_{\lambda}(U_{i}(x))$.
If $X=D(c,\delta,\varepsilon)$ is an annulus, then apply the same arguments as in [11, Lemma 3.7]. We may split of any number of intervals, since we already know how to deal with these. So let $D=D(c,\delta,\varepsilon)$ be one of the finitely many resulting annuli. The arguments of [11, Lemma 3.7] show that
@@ 3.5. Pulling back from the real closure
We use induction on $n=\deg(p)$. If $n=1$ it is clear, since then we can simply take $C=\{\alpha_{1}\}$. So we assume that $n>1$ and that the statement is proven for smaller $n$. By induction, let $C$ be a finite set for the claim applied to $p^{\prime}$, and also such that $\operatorname{rv}(p(x))$ is constant on balls $1$-next to $C$, which is possible by [19, Prop. 3.6]. This $C$ is ${\mathcal{L}}_{\mathrm{val}}(A)$-definable, and we claim that it suffices. So let $x,y$ be in the same ball $1$-next to $C$, and assume that the claim is false. Then there exists some root $\alpha$ of $p$ such that $\operatorname{rv}(x-\alpha)\neq\operatorname{rv}(x-\alpha)$. Let $B\subset K^{\mathrm{rcl}}$ be the closed valuative ball around $x$ of radius $v(x-y)$, then $\alpha$ is an element of $B$. If $B$ would contain another root $\alpha^{\prime}\in K^{\mathrm{rcl}}$ of $p$, then by Rolle’s theorem it would also contain a root of $p^{\prime}$, since $K^{\mathrm{rcl}}$ is real closed and $B$ is convex. But this contradicts the construction of $C$, and so the only root of $p$ in $B$ is $\alpha$. Hence if $\alpha^{\prime}$ is another root of $p$, then $\operatorname{rv}(x-\alpha^{\prime})=\operatorname{rv}(y-\alpha^{\prime})$. But if we now compute
@@ 4.1. Eliminating to $\mathrm{RV}_{\lambda}$
The proof of h-minimality is based on relative quantifier elimination results in a larger many-sorted language. First we apply the strategy of Denef–van den Dries [14] to reduce to the case where the variables over which we quantify occur only polynomially. Then we apply the relative quantifier elimination of result of Flenner [19] for the pure valued field structure. To go from relative quantifier elimination to $\omega$-h-minimality, we follow the method from [6, Sec. 6].
@@
We call this the partial addition. We recall a first relative quantifier elimination result due to Flenner (see also [3] for an earlier result in a related language).
4.1.1 Theorem ([19, Prop. 4.3]).
@@
where $z=(z_{1},\ldots,z_{m})$ are valued field variables, the $f_{i}$ are either in $B_{m+n,\alpha}$ or a polynomial over $K$, and no field division occurs. Let $(z_{0},y_{0})\in[-1,1]^{m+n}$. Using the Strong Noetherian property from [8, Theorem 3.2.2], there exist an integer $d$ and a real number $\beta(z_{0},y_{0})>0$, unit elements $u_{ij}\in B^{\circ}_{m+n,\beta}$, elements $h_{ij}\in B_{n,\beta}$ and a subset $J_{i}$ of $\{0,...,d\}^{m}$ such that for every $i=1,\ldots,s$,
@@
for finitely many points $(z_{0},y_{0})\in[-1,1]_{{\mathbb{R}}}^{m+n}$. Without loss of generality, we can assume that $(z_{0},y_{0})=(0,0)$. Now, by a coordinate transformation and Weierstrass division similar to [11, Proof of Theorem 2.14] and [14, Lemma 4.13 and Proof 4.14], we can shrink $\beta$ if needed and write $f_{i}(z,y)$ as
@@
for some fixed $j$, $c\in[-1,1]_{{\mathbb{R}}}$, where $v_{i}$ is a unit and $g_{i}$ polynomial in the last variable $z_{m}$. Now, if $v\in B_{n^{\prime},\alpha}^{\circ}$ is a unit, then by construction of our language, there exists a definable map $\overline{v}$ on $\mathrm{RV}_{\lambda}^{n^{\prime}}$ such that $\operatorname{rv}_{\lambda}(v(x))=\overline{v}(\operatorname{rv}_{\lambda}(x))$. Hence $\operatorname{rv}_{\lambda}(f_{i}(z,y))$ depends only on $\operatorname{rv}_{\lambda}(z,y),\operatorname{rv}_{\lambda}(h_{i\ell}(y)/h_{j\ell}(y)-c)$ and $\operatorname{rv}_{\lambda}(g_{i}(y,h_{i\ell}(y)/h_{j\ell}(y)-c,z))$. In other words, the variable $z_{m}$ only occurs polynomially in our formula (with some parameters depending on $y$). If $\lambda=1$, we can then conclude from Theorem 4.1.1 to eliminate the quantifier over $z_{m}$ and iterate this procedure. If $\lambda\neq 1$, then the result follows from [6, Prop. 6.2.3]. ∎